Using the method Laplace of separation of variables, solve du /dt=d²u/dx², u(0,t)=0, u(4,t)=0, u(x,0)=sin3x


1
Expert's answer
2021-05-18T18:13:18-0400

the problem is posed incorrectly since the initial conditions and boundary conditions contradict each other:


"u(x,0)=sin(3x) \space and \space u(4,t)=0, but \space sin(3\cdot4) \not=0"

Suppose


"u(x,0)=sin(3\pi x)"

We apply the Laplace transform to the function u(x,t) in t, considering x as a parameter


"L\{ u(x,t) \}=U(x,s)"

"L\Big\{ \frac {d u(x,t)}{dt} \Big\}=sU(x,s)-u(x,0)=sU(x,s)-sin(3\pi x)"

"L\Big\{ \frac {d^2 u(x,t)}{dx^2} \Big\}=U_{xx}(x,s)"

Applying the Laplace transform to the equation, we have


"U_{xx}(x,s)=sU(x,s)-sin(3\pi x)"

The general solution can be written as


"U(x,s)=U_h(x,s)+U_p(x,s)"

the general solution of the homogeneous equation


"U_h(x,s)=C_1 e^{x \sqrt s}+C_2 e^{-x \sqrt s}"

the particular solution of the non-homogeneous equation


"U_p(x,s)=C_3 cos(3\pi x)+C_4 sin(3\pi x)"

"\frac {d^2 U_p(x,s)}{dx^2}=-(3\pi)^2 U_p(x,s)"

substitute the particular solution into the equation


"-(3\pi)^2 U_p(x,s)=sU_p(x,s)-sin(3\pi x)"

"(-(3\pi)^2-s)(C_3 cos(3\pi x)+C_4 sin(3\pi x))=-sin(3\pi x)"

"(-(3\pi)^2-s)C_3=0\space \rightarrow \space C_3=0"

"(-(3\pi)^2-s)C_4=-1\space \rightarrow \space C_4=\frac{1}{s+9\pi^2}"

the general solution


"U(x,s)=C_1 e^{x \sqrt s}+C_2 e^{-x \sqrt s}+\frac{sin(3\pi x)}{s+9\pi^2}"

from the boundary condition


"u(0,t)=0 \rightarrow U(0,s)=0=C_1 +C_2+0"

"u(4,t)=0 \rightarrow U(4,s)=0=C_1 e^{4 \sqrt s}+C_2 e^{-4 \sqrt s}+0"

then


"C_1 =C_2 =0"

So, the solution is


"U(x,s)=\frac{sin(3\pi x)}{s+9\pi^2}"

apply the inverse Laplace transform


"u(x,t)=L^{-1} \Big\{ \frac{sin(3\pi x)}{s+9\pi^2} \Big\}=sin(3\pi x)L^{-1} \Big\{ \frac{1}{s+9\pi^2} \Big\}"

"u(x,t)=e^{-9\pi^2t}sin(3\pi x), \space t\ge0"

Answer:"\space u(x,t)=e^{-9\pi^2t}sin(3\pi x), \space t\ge0, \space 0\le x\le 4"